3.4Suppose that the real GDP increase by R5,000 billion when government expenditure on the construction of new roads increase by R1,500 billion. What is the value of the marginal propensity to consume?
3.5Assuming that the central government decides to cut taxes by R100 billion to stimulate the economy. The relevant marginal propensity to consume is 0.6 (60 percent). What will be the impact of such fiscal policy on equilibrium GDP?
1
Expert's answer
2020-04-30T10:05:49-0400

3.4 Suppose that the real GDP increase by R5,000 billion when government expenditure on the construction of new roads increase by R1,500 billion. What is the value of the marginal propensity to consume?

"MPC = ( \u25b3C)\/(\u25b3y)"

"= (1500\/5000)= 0.3"

3.5

Marginal Propensity to consume + marginal propensity to save = 1 (enclosed system)

"MPC + MPS = 1"

"0.6 +MPS =1"

"MPS = 1-0.6= 0.4"

"Tax multiplier = (-MPC)\/MPS)"

"= (-0.6)\/0.4)"

"-1.5"

Final impact on GDP=multiplier × autonomous change in taxation

"= -1.5 \u00d7 -R100 = R150"

Tax cut by R100 had an impact of increasing GDP by R150 billion